OpenStudy (anonymous):

What is the solution set of this system of equations? y = x2 − 3x − 4 x = y + 8 a.{(-1, 0), (4, 0)} B.{(8, 0), (0, 8)} C.{(0, 0)} D.{(2, -6)} E.there is no solution

OpenStudy (anonymous):

Plug in your second equation to your first y = x2 − 3x − 4 x = y + 8 So by plugging it in we can get y=(y+8)^2-3(y+8)-4 <--- solve y here

OpenStudy (muzzack):

E lol